VEGAS Pulsar Project Meeting: 2016 January 11th 2:00 - 3:00pm ET

Room / Connection Details

  • GB-137 / Soc 280
  • 434-817-6443

Housekeeping Details

  • Main wiki page
  • gbsapp Mailing List
  • Location and regular date and time of meeting - agreed weekly on Mondays; GB-137 / Soc-280; 2-3pm ET
  • ETK codes:
    • Electronics: 432510.GB3015
    • Software: 893252.GB3015
    • Science: 432540.GB3015

Resource Allocations

LYNCH, Ryan 0.15


  • Randy
  • Justin
  • Ray
  • Richard
  • Paul
  • Marty


BOF file status updates

  • c1500x4096 mode
    • Recompiled GPU transpose code
      • DONE Complete
    • nVidia profiling
      • Likely not complete
      • BLI found using packet-sockets solves this problem and found them highly efficient. That code is done and available.
    • guppi_daq_server version comparisons
      • Justin to test for performance differences
    • Is this the SS version?
  • c1500x0512 mode
    • Build status
      • Built successfully
    • Scans failing during long integrations
      • Update status shared memory code to use different data type to accurately capture packet ID
        • only for displaying packet, does not solve failure
      • Cause of failure is not yet determined
        • difficult to reproduce
        • Needs a lead
          • Justin to talk to Ryan
    • Astronomical verification
      • Used successfully during Breakthrough Listen tests (though with some packet loss in the RAW mode)
        • Packet loss problem now solved using the BLI packet-socket code
    • Is this the SS version?
  • c1500x2048 mode
    • inability to build off of c1500x4096
      • still have not determined cause
      • looking at block-level designs to localize issues
    • Is this the SS version?
  • Fix for nchan calculation in coherent modes
    • verified - ready to roll into other modes
  • Bandpass issues in i0800x0064, 0128, 8192
    • Spent some time testing 64 channel mode
    • Tone shows up in the correct bins, but generates spurs if injected at Fs/8 or nearby
      • Spurs move with the test tone and decrease with amplitude if you move away from Fs/8
      • May be problem with MMCM or OGP calibration
        • Update on calibration
    • A suggestion was made to build a BOF with digital noise source to test
      • tough for 8192, others should be ok
        • test w/ 64 chl mode since testing already done by Jason
      • work needs to prioritized against other work
  • Duplicated signal in i0800x8192
    • Unknown by attendees if any work has been done
    • Related to above issue
  • Low bandwidth modes
    • New filter coefficients
    • Spectrum output instead of time series
    • Implement PFB code
    • Could single GPU keep up with coherent dedispersion?
    • Reviewed John's email suggestion with Paul - more detailed query with John recommended
      • Is there an advantage to FFTs in FPGA? (none identified)
      • GPU (S/W) vs. FPGA (H/W)
        • GPU likely preferable if it can handle it
          • requires a benchmarking test
          • existing FPGA filter work could be affected
        • S/W CUDA-savvy resource constrained
          • Justin will look at it
      • How about running a higher bandwidth modes and ignoring some of the spectra?
        • limits on number of channels likely an issue


  • Configtool changes
    • Status of rebasing on current release version
      • after next M&C Release
      • dual backend code added but not system tested
        • Justin to test on next maintenance day
  • GBT Status shared memory parameters in PSRFITS headers
    • Done and seems to be working well
  • Switching signal generation
    • Cal successfully controlled by VEGAS
    • Actual swper still not matching requested swper
      • Status of calculation of actual swper
        • no progress/solution not clear
  • Matrix HPC
    • built compliance for all threads in Guppi Dac Server
      • some require additional work

GBT testing

  • No updates

Wideband AP

  • Status of new filters
    • on order, delayed by ITAR issues, now resolved
      • Due in ~ 3wks
      • Populating circuit boards with other parts
  • Any other updates

VEGAS/DIBAS-style backend for FAST

  • John and Richard met with Tony B and Karen on Dec 22
    • Updates from that meeting
  • No response to email asking FAST to send request to T. Beasley

Multicasting for spectral line and pulsar modes

  • Modifications to ROACH configuration and net task must bind to a multicast group instead of a specific host
    • All the 10-gig addresses on the ROACH would be multicast groups, and config will be told which to bind to
    • Some changes to net task code (in VEGAS HPC and guppi_daq code) to initialize sockets

Resource Allocations for next quarter

  • See above

Date / Time of Next Meeting

  • January 18
-- RyanLynch - 2015-12-30

-- MartyBloss - 2016-01-12
Topic revision: r3 - 2016-01-12, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback